# Advancing Averaged Primer Vector Theory with Bang-Bang Control and Eclipsing

### v1

@@ 6 Conclusion
An improved model is introduced for averaged dynamics and sensitivities for the minimum fuel problem that combines advances from several past works with new developments in the eclipsing constraint, the switching function, an expanded state vector, and variational equations. A singularity in the averaged eclipsing constraint is removed, allowing a variable step integrator to propagate through small eclipsing arcs without stopping, significantly slowing down, or accumulating error. The roots of the switching function are shown to correspond to a sixth order polynomial in the averaging context, contributing to faster root-finding and thus faster propagation of multi-arc averaged dynamics. This result implies a maximum of three thrusting arcs per revolution in averaged dynamics with deeper implications for the upper bound on number of thrusts per revolution for optimal transfers in general. The variational equations for averaged dynamics are used to calculate an accurate and consistent STM, enabling a smoother optimization problem. These new developments enhance the applicability of the averaged model to optimizing low-thrust, many revolution trajectories with bang-bang control and a discontinuous eclipsing constraint. In practice, the averaged model introduced here can be the final step of a continuation process used in solving the BVP. The initial and intermediary steps of this continuation process can use quadratic cost functions or smooth Heaviside approximations in the averaged model.
Appendix A
The coefficients for the switching polynomial are shown in Table 4 with parentheses used for algebraic groupings and the initial conditions for the example trajectories are shown in Table 5. The matrices A and B for the equations of motion of modified equinoctial orbital elements are:

### v3

@@ 6 Conclusion
An improved model is introduced for averaged dynamics and sensitivities for the minimum fuel problem that combines advances from several past works with new developments in the eclipsing constraint, the switching function, an expanded state vector, and variational equations. A singularity in the averaged eclipsing constraint is removed, allowing a variable step integrator to propagate through small eclipsing arcs without stopping, significantly slowing down, or accumulating error. The roots of the switching function are shown to correspond to a sixth-order polynomial in the averaging context, contributing to faster root-finding and thus faster propagation of multi-arc averaged dynamics. This result implies a maximum of three thrusting arcs per revolution in averaged dynamics with deeper implications for the upper bound on number of thrusts per revolution for optimal transfers in general. The variational equations for averaged dynamics are used to calculate an accurate and consistent STM, enabling a smoother optimization problem. These new developments enhance the applicability of the averaged model to optimizing low-thrust, many revolution trajectories with bang-bang control and a discontinuous eclipsing constraint. In practice, the averaged model introduced here can be the final step of a continuation process used in solving the boundary value problem. The initial and intermediary steps of this continuation process can use quadratic cost functions or smooth Heaviside approximations in the averaged model.
Appendix
The coefficients for the switching polynomial are shown in Table Appendix with parentheses used for algebraic groupings and the initial conditions for the example trajectories are shown in Table Appendix. The matrices A and B for the equations of motion of modified equinoctial orbital elements are:
@@
Funding Sources
This work was supported in part by a NASA Space Technology Graduate Research Opportunity, grant number 80NSSC22K1209.
Acknowledgments
The authors thank Steve McCarty for technical motivation and advice.